OpenStudy (anonymous):

Medal, Fan, Fan Testimonial PLZ Help A crater shaped like a cone has a height of 24 yd and a diameter of 71 yd. What is the volume of the crater? Use 3.14 to approximate pi and round your answer to one decimal place. Answer is in >>>yd3<<<

OpenStudy (anonymous):

@Abhisar @AnswerMyQuestions @bohotness @DarkBlueChocobo @freckles @hhelpplzzzz @Luigi0210 @z

OpenStudy (anonymous):

@Preetha @dakari22

OpenStudy (anonymous):

The formula is \(\dfrac{1}{3}\) x \(\LARGE\pi\) x \(\Large{r^2}\) x \(h\)

OpenStudy (anonymous):

So after we plug in the values, we have \(\dfrac{1}{3}\) x 3.14 x 35.5^2 x 24, so can you figure that out? @MAli13chineta

OpenStudy (anonymous):

did u figure it out cuz i need the answer